Suppose Rural wage = $2.40 per day, Urban modern wage = $4.50 per day, Urban informal sector income = $1.50 per day. Suppose there is a 35% probability of getting a modern job. Based on the above information, applying the Harris-Todaro model, answer the following questions:

i) Will there be an incentive for a worker to migrate from rural to urban sector?

ii) At what probability of getting a modern job, will the worker be indifferent between choosing a rural and an urban job, given the above wages?
Answer #1

Ans.1- Expected wage in urban sector = 0.35(4.5) + 0.65(1.5) = 2.55

Since expected wage in urban sector is higher than the rural sector so there will be an incentive to migrate from rural to urban sector.

Ans.2- Worker will be indifferent if

Expected wage in urban sector= Rural wage = 2.4

Let p be the probability to get a modern job

p(4.5) + (1-p)(1.5) = 2.4

3p+ 1.5 = 2.4

3p= 0.9

p= 0.3

So,at 30% probability worker will be indifferent .